OpenStudy (anonymous):

3a^4b^-2c^3? A)81a^4c^3/b^2 B) 81a^4/b^2c^3 C)3a^4/b^2c^3 D) 3a^4c^3/b^2

OpenStudy (anonymous):

@lackofname

OpenStudy (anonymous):

i have no idea... i dont even know what to do ...

OpenStudy (anonymous):

just go with c lol

OpenStudy (anonymous):

No you have solve it

OpenStudy (anonymous):

(3)^4*(4)^-2*(2)^3

OpenStudy (phebe):

C

OpenStudy (anonymous):

3*a^4*b^-2*c^3 = (3a^4*c^3)/(b^2) b^-2 = 1/b^2